OpenStudy (anonymous):

The first outcome is one of A and B the second outcome in each case is one of a and b. Use the following probabilities: Pr[A] = [ 4/5] Pr[B] = [ 1/5] Pr[a | A] = [ 7/9] Pr[b | A] = [ 2/9] Pr[a | B] = [ 4/7] Pr[b | B] = [ 3/7] Pr[b] = [ 83/315 ] What is Pr[B | b]?

OpenStudy (anonymous):

I thought it would be Pr[B]/Pr[b], but it says that is wrong. So not 63/83

OpenStudy (anonymous):

@DanielxAK @satellite73

OpenStudy (anonymous):

Pr[B|b] = Pr[Bb]/Pr[b]. You don't know what Pr[Bb] is, since they don't give it to you. Can you think of another way to write Pr[Bb] in terms of what you know?

OpenStudy (anonymous):

Pr[Bb] is (3/35)... I found Pr[Ba] to be 4/35 using Pr[a | B] = [ 4/7], so 4/35 + Bb=1/5, so Bb= (3/35)

OpenStudy (anonymous):

Okay, so you should be able to solve it from there, correct?

OpenStudy (anonymous):

That would be (1/5)/(3/35), but that comes out to 7/3, which it also says is wrong!

OpenStudy (anonymous):

P[B|b] = P[Bb]/P[b] Check what numbers you're plugging into your equation.

OpenStudy (anonymous):

That worked! But why is it Pr[Bb]/Pr[b] and not Pr[B]/Pr[b]?

OpenStudy (anonymous):

That's the definition. Given 2 events, A and B, you have: P[A|B] = P[AB]/P[B] P[AB] = P[A] only if B was a subset of A.

OpenStudy (anonymous):

That explains a lot. Thank you!!